A 30g sample of ore contains 25g of haematite, Fe2O3. What mass of iron could be extracted?
1
Expert's answer
2020-05-03T15:14:40-0400

As you can see from the molecular formula, there are 2 atoms of iron in one molecule of haematite Fe2O3. Therefore the number of the moles of hematite and the number of the moles of iron that can be extracted relate as:

"n(Fe_2O_3) = n(Fe)\/2" .

Let's calculate the number of the moles of haematite. Its molar mass "M" is 159.69 g/mol:

"n(Fe_2O_3) = \frac{m}{M} = \frac{25}{159.69} = 0.157" mol.

Then, the number of the moles of iron is:

"n(Fe) = 2\u00b70.157 = 0.313" mol.

The molar mass of iron is 55.84 g/mol. The mass of the iron that can be extracted:

"m(Fe) = n\u00b7M = 0.313\u00b755.84 =17.5" g

Answer: A 30g sample of ore contains 25g of haematite, Fe2O3. The mass of iron could be extracted is 17.5 g.